# Local logistic regression for dimension reduction in classification

### v1

@@ 3.3 Real data analyses
  • The Hill-Valley (HV) dataset [^2] https://archive.ics.uci.edu/dataset/166/hill+valley. Each data point is made of 100 real numbers $\bm{x}_{i}=(x_{i,j})_{1\leq j\leq 100}$ which, when plotted in order, create a curve in the two-dimensional plane that features a hill (a “bump” in the curve) or a valley (a “dip” in the curve). The data consists of the $n=1212$ pairs $(Y_{i},\bm{x}_{i})\in\{0,1\}\times\mathbb{R}^{100}$, where $Y_{i}=0$ if the $i$th curve features a valley and 1 if it features a hill.
@@
Each dataset is divided at random into a training set and a testing set, approximately made of $70\%$ and $30\%$ of the original data, respectively. Table 1 compares the performance of the complete workflow (dimension reduction and selection and then classification) using the full space of covariates, the covariates projected on the dimension reduction subspace provided by the proposed LLO$(\lambda>0)$ method, and its version obtained using the non-penalized version LLO$(\lambda=0)$, paired with either the knn or the RandomForest classifier, when applied to the testing set. The classification procedure using LLO$(\lambda>0)$ generally has a comparable or lower misclassification risk, a comparable or higher AUC (this can also be seen by comparing the ROC curves of each classification procedure, see Figure S.6) with comparable or lower computing time with respect to the non-penalized version, and always improves substantially upon the classifier not featuring dimension reduction. It is interesting to note that, as expected, the cross-validation procedure for the dimension of the reduction subspace involving the estimation of the matrix $M$ via LLO$(\lambda>0)$ tends to select fewer components than its analogue using LLO$(\lambda=0)$, with comparable or higher accuracy, as shown in the top panels in Figure 5 for the nearest neighbor classifier (see Figure S.5 for similar results with the Random Forest classifier). It is noted that in the WBDC real data analysis, all the eigenvalues of the empirical outer product $\widehat{M}$ were found to be 0 from dimension $d=15$ and $d=22$ onwards when using the LLO($\lambda>0$) and LLO($\lambda=0$) method, respectively.
Table 1: Misclassification risk, the area under the ROC curve, and computing time for each classifier applied to the three real datasets. In each case, the prediction exercise is carried out on the selected testing set.
@@
Miscl. risk AUC Est. time (secs) Miscl. risk AUC Est. time (secs)
@@
We also compare these classifiers to those obtained following dimension reduction via the SAVE, PHD, and POTD methods. The misclassification risk is compared for estimated central subspaces of dimension $d\in\{2,3,4,5,10\}$. It can be seen in the bottom panels of Figure 5 that the classifier obtained after dimension reduction through the proposed LLO$(\lambda>0)$ approach is consistently superior to its competitors. The classifier based on LLO$(\lambda=0)$ dimension reduction performs worse but its performance gets closer to that of the LLO$(\lambda>0)$-based classifier as the dimension of the reduction subspace increases. Similar conclusions can be reached by considering the ROC curves of the different classifiers; see Figures S.7S.9.
@@ Acknowledgments
“Local logistic regression for dimension reduction in classification”
@@
S.1.2 Functional weak convergence of certain nearest-neighbor empirical processes (proof of Theorem 1)
We first recall the following lemma on the convergence of the $k$-NN bandwidth; see Lemma 1 in Portier (2021).
@@
Suppose that (A1) is fulfilled. Let $\Psi_{n}:\{0,1\}\times\mathbb{R}^{p}\to\mathbb{R}^{q}$ be a sequence of measurable vector-valued functions and suppose that there is a positive integer $n_{0}$ such that
$\Psi_{\infty}:=\sup_{n\geq n_{0}}\sup_{z\in B(x,(3/2)^{1/p}\tau_{n,k}(x))}\max(\|\Psi_{n}(0,z)\|_{2},\|\Psi_{n}(1,z)\|_{2})<\infty.$
@@
For two functions $\underline{f},\overline{f}\in L_{2}(Q)$, the bracket $[\underline{f},\overline{f}]$ is the set of all functions $g$ in $L_{2}(Q)$ such that $\underline{f}\leq g\leq\overline{f}$ on $S$. A bracket $[\underline{f},\overline{f}]$ such that $\|\underline{f}-\overline{f}\|_{L_{2}(Q)}\leq\varepsilon$ is called an $\varepsilon$-bracket, and for any $\mathcal{G}\subset L_{2}(Q)$, the $\varepsilon$-bracketing number, denoted by $\mathcal{N}_{[\,]}(\mathcal{G},L_{2}(Q),\varepsilon)$, is defined as the smallest number of $\varepsilon$-brackets needed to cover $\mathcal{G}$. Finally, call an envelope function for $\mathcal{G}$ any function $G:S\mapsto\mathbb{R}$ such that $|g|\leq G$ on $S$ for any $g\in\mathcal{G}$.
@@
with
$\Psi_{n}(y,z)=\varphi(z)\begin{pmatrix}1\\ \frac{z-x}{\tau_{n,k}(x)}\end{pmatrix}\begin{pmatrix}1\\ \frac{z-x}{\tau_{n,k}(x)}\end{pmatrix}^{T}.$
@@
$\sup_{n\geq n_{0}}\sup_{z\in B(x,(3/2)^{1/p}\tau_{n,k}(x))}\max(\|\Psi_{n}(0,z)\|_{2},\|\Psi_{n}(1,z)\|_{2})\leq|\varphi|_{\infty}\left(1+\left(\frac{3}{2}\right)^{2/p}\right).$
@@
$\Psi_{n}(y,z)=\left\{\left(\frac{z-x}{\tau_{n,k}(x)}\right)^{T}M\left(\frac{z-x}{\tau_{n,k}(x)}\right)\right\}\frac{z-x}{\tau_{n,k}(x)}$
@@
with
@@
$+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(cf_{X}(x)V_{p})^{1/p}}\left(\sum_{j=1}^{p}\mathrm{sgn}(\nabla\ell_{j}(x))u_{j}\mathbbm{1}_{\{\nabla\ell_{j}(x)\neq 0\}}+|u_{j}|\mathbbm{1}_{\{\nabla\ell_{j}(x)=0\}}\right).$
@@
and $\mbox{and }A_{n,2}(u)$ $\stackrel{{\scriptstyle\mathbb{P}}}{{\longrightarrow}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(cf_{X}(x)V_{p})^{1/p}}\left\{\sum_{j=1}^{p}\mathrm{sgn}(\nabla\ell_{j}(x))u_{j}\mathbbm{1}_{\{\nabla\ell_{j}(x)\neq 0\}}+|u_{j}|\mathbbm{1}_{\{\nabla\ell_{j}(x)=0\}}\right\}.$
@@
We apply (S.7) with $F:u\mapsto L_{n}(z+D_{n}u)$ where $z=(\ell(x),\nabla\ell(x)^{T})^{T}$. We know that the gradient of $F$ at $0_{p+1}$ is $S_{n}(x)$ and its Hessian matrix at $0_{p+1}$ is $H_{n}(x)$; more generally
@@
$\|H_{F}(u)-H_{F}(0_{p+1})\|_{2}\leq\frac{1}{\sqrt{k}}\times\frac{1}{4}\|u\|_{2}\times\frac{1}{k}\sum_{i\in N_{k}(x)}\left\|\begin{pmatrix}1\\ \frac{X_{i}-x}{\tau_{n,k}(x)}\end{pmatrix}\right\|_{2}^{3}.$
@@
Figure S.6: Real data analysis – ROC curve of the nearest neighbor classifier (left) and random forest classifier (right) with no dimension reduction (black curve), dimension reduction following the LLO$(\lambda=0)$ procedure (red curve) and dimension reduction following the LLO$(\lambda>0)$ procedure (blue curve). In each case, the prediction exercise is carried out on the selected testing set.

### v2

@@ 3.3 Real data analyses
  • The Hill-Valley (HV) dataset [^2] https://archive.ics.uci.edu/dataset/166/hill+valley. Each data point is made of 100 real numbers $\bm{x}_{i}=(x_{i,j})_{1\leq j\leq 100}$ which create a curve in the two-dimensional plane that features a hill (a “bump” in the curve) or a valley (a “dip” in the curve). The data consists of the $n=1212$ pairs $(Y_{i},\bm{x}_{i})\in\{0,1\}\times\mathbb{R}^{100}$, where $Y_{i}=1$ if and only if the curve features a hill.
@@
Figure 4: Simulation study – Left panel: Dimension selection through Algorithm 2, where the number indicated above the curve gives the number of times the dimension selected in the (absolute or relative) majority of cases was chosen. Right panel: Misclassification risk of the nearest-neighbor classifier with, from left to right, $d=p$ (red bar), the covariates projected on the correct population central subspace (blue bar), the central subspace estimated using the non-penalized LLO($\lambda=0$) method under correct specification of the dimension (green bar) and with the dimension estimated by cross-validation (purple bar), and the central subspace estimated using the penalized LLO($\lambda>0$) method under correct specification of the dimension (orange bar) and with the dimension estimated by cross-validation (yellow bar). Both panels are produced using $N=1000$ independent replications of a sample of size $n=1000$ and dimensions $p\in\{8,16,32,64\}$ of the full covariate space are considered.
Each dataset is divided at random into a training set and a testing set, approximately made of $70\%$ and $30\%$ of the original data, respectively. Table 1 compares the performance of the complete workflow (dimension reduction and selection and then classification) using the full space of covariates, the covariates projected on the dimension reduction subspace provided by the proposed LLO$(\lambda>0)$ method, and its version obtained using the non-penalized version LLO$(\lambda=0)$, paired with either the knn or the RandomForest classifier, when applied to the testing set. The classification procedure using LLO$(\lambda>0)$ generally has a comparable or lower misclassification risk, a comparable or higher AUC (this can also be seen by comparing the ROC curves of each classification procedure, see Figure S.6) with comparable or lower computing time with respect to the non-penalized version, and always improves substantially upon the classifier not featuring dimension reduction. We note that, as expected, the cross-validation procedure for the dimension of the reduction subspace involving the estimation of the matrix $M$ via LLO$(\lambda>0)$ tends to select fewer components than its analogue using LLO$(\lambda=0)$, with comparable or higher accuracy, as shown in the top panels in Figure 5 for the nearest neighbor classifier (see Figure S.5 for similar results with the Random Forest classifier). It is noted that in the WBDC real data analysis, all the eigenvalues of the empirical outer product $\widehat{M}$ were found to be 0 from dimension $d=15$ and $d=22$ onwards when using the LLO($\lambda>0$) and LLO($\lambda=0$) method, respectively.
Table 1: Misclassification risk, the area under the ROC curve, and computing time (in seconds) for each classifier applied to the three real datasets. In each case, the prediction exercise is carried out on the selected testing set.
@@
Miscl. risk AUC Est. time Miscl. risk AUC Est. time
@@
We also compare these classifiers to those obtained following dimension reduction via the SAVE, PHD, and POTD methods. The misclassification risk is compared for estimated central subspaces of dimension $d\in\{2,3,4,5,10\}$. It can be seen in the bottom panels of Figure 5 that the classifier obtained after dimension reduction through the proposed LLO$(\lambda>0)$ approach is consistently superior to its competitors. The classifier based on LLO$(\lambda=0)$ dimension reduction performs worse but its performance gets closer to that of the LLO$(\lambda>0)$-based classifier as the dimension of the reduction subspace increases. Similar conclusions can be reached by considering ROC curves; see Figures S.7S.9.
(a) HV – Dimension selection
(b) HV – Dimension selection
(c) MPE – Dimension selection
(d) HV – Comparison between methods
(e) MPE – Comparison between methods
(f) WDBC – Comparison between methods
Figure 5: Real data analysis - Top panels: Dimension selection through cross-validation for LLO($\lambda=0$) and LLO($\lambda>0$), using the nearest neighbor classifier. Bottom panels: Estimated misclassification risk related to the nearest neighbor classifier on the set of projected covariates on the estimated subspace produced by each method for a dimension $d$ of the central subspace in $\{2,3,4,5,10\}$, with 95% asymptotic Gaussian Wald-type confidence intervals (the sample size used is the size of the testing set).
@@ Acknowledgments
“Logistic lasso regression with nearest neighbors for gradient-based dimension reduction”
@@
S.1.2 Functional weak convergence of certain nearest-neighbor empirical processes (Theorem 4 and its proof)
The goal of this section is to prove the following central limit theorem for nearest-neighbor estimators, which, besides being the crucial result behind our asymptotic theory, is of independent interest.
Theorem 4 (Central limit theorem for nearest-neighbor estimators).
Let $E$ be a nonempty and finite set. Assume that the data is made of independent copies $(Y_{i},X_{i})$ of the random pair $(Y,X)\in E\times\mathbb{R}^{p}$ and that (A1) is fulfilled. Assume that $k:=k_{n}\to\infty$ is such that $k/n\to 0$. Let $\Psi_{n}:E\times\mathbb{R}^{p}\to\mathbb{R}^{q}$ be a sequence of measurable vector-valued functions and suppose that there is a positive integer $n_{0}$ such that
$\sup_{n\geq n_{0}}\sup_{z\in B(x,(3/2)^{1/p}\tau_{n,k}(x))}\max_{y\in E}\|\Psi_{n}(y,z)\|_{2}<\infty.$
  1. (i)
    Then $Z_{n}(\widehat{\tau}_{n,k}(x))=O_{\mathbb{P}}(1)$.
  2. (ii)
    then
To prove this result, we first recall the following lemma on the convergence of the $k$-NN bandwidth; see Lemma 1 in Portier (2021).
@@
Let $E$ be a nonempty and finite set. Assume that the data is made of independent copies $(Y_{i},X_{i})$ of the random pair $(Y,X)\in E\times\mathbb{R}^{p}$ and that (A1) is fulfilled. Let $\Psi_{n}:E\times\mathbb{R}^{p}\to\mathbb{R}^{q}$ be a sequence of measurable vector-valued functions and suppose that there is a positive integer $n_{0}$ such that
$\Psi_{\infty}:=\sup_{n\geq n_{0}}\sup_{z\in B(x,(3/2)^{1/p}\tau_{n,k}(x))}\max_{y\in E}\|\Psi_{n}(y,z)\|_{2}<\infty.$
@@
For two functions $\underline{f},\overline{f}\in L_{2}(Q)$, the bracket $[\underline{f},\overline{f}]$ is the set of all functions $g$ in $L_{2}(Q)$ such that $\underline{f}\leq g\leq\overline{f}$ on $S$. A bracket $[\underline{f},\overline{f}]$ such that $\|\underline{f}-\overline{f}\|_{L_{2}(Q)}\leq\varepsilon$ is called an $\varepsilon$-bracket, and for any $\mathcal{G}\subset L_{2}(Q)$, the $\varepsilon$-bracketing number, denoted by $\mathcal{N}_{[\,]}(\mathcal{G},L_{2}(Q),\varepsilon)$, is defined as the smallest number of $\varepsilon$-brackets needed to cover $\mathcal{G}$. Finally, call an envelope function for $\mathcal{G}$ any function $G:S\to\mathbb{R}$ such that $|g|\leq G$ on $S$ for any $g\in\mathcal{G}$.
@@
with $E=\{0\}$, $Y_{i}=0$ for all $i$ and
$\Psi_{n}(y,z)\equiv\Psi_{n}(z)=\varphi(z)\begin{pmatrix}1\\ \frac{z-x}{\tau_{n,k}(x)}\end{pmatrix}\begin{pmatrix}1\\ \frac{z-x}{\tau_{n,k}(x)}\end{pmatrix}^{T}.$
@@
$\sup_{n\geq n_{0}}\sup_{z\in B(x,(3/2)^{1/p}\tau_{n,k}(x))}\|\Psi_{n}(z)\|_{2}\leq|\varphi|_{\infty}\left(1+\left(\frac{3}{2}\right)^{2/p}\right).$
@@
$\Psi_{n}(y,z)\equiv\Psi_{n}(z)=\left\{\left(\frac{z-x}{\tau_{n,k}(x)}\right)^{T}M\left(\frac{z-x}{\tau_{n,k}(x)}\right)\right\}\frac{z-x}{\tau_{n,k}(x)}$
@@
with $\Psi_{n}:E\times\mathbb{R}^{p}=\{0,1\}\times\mathbb{R}^{p}\to\mathbb{R}^{p+1}$ defined by
@@
$+(cf_{X}(x)V_{p})^{1/p}\left(\sum_{j=1}^{p}\mathrm{sgn}(\nabla\ell_{j}(x))u_{j}\mathbbm{1}_{\{\nabla\ell_{j}(x)\neq 0\}}+|u_{j}|\mathbbm{1}_{\{\nabla\ell_{j}(x)=0\}}\right).$
@@
and $\mbox{and }A_{n,2}(u)$ $\stackrel{{\scriptstyle\mathbb{P}}}{{\longrightarrow}}(cf_{X}(x)V_{p})^{1/p}\left\{\sum_{j=1}^{p}\mathrm{sgn}(\nabla\ell_{j}(x))u_{j}\mathbbm{1}_{\{\nabla\ell_{j}(x)\neq 0\}}+|u_{j}|\mathbbm{1}_{\{\nabla\ell_{j}(x)=0\}}\right\}.$
@@
We apply (S.7) with $F:u\mapsto L_{n}(z+D_{n}u)$ where $z=(\ell(x),\nabla\ell(x)^{T})^{T}$. We know that the gradient of $F$ at $0=0_{p+1}$ is $S_{n}(x)$ and its Hessian matrix at 00 is $H_{n}(x)$; more generally
@@
$\|H_{F}(u)-H_{F}(0)\|_{2}\leq\frac{1}{\sqrt{k}}\times\frac{1}{4}\|u\|_{2}\times\frac{1}{k}\sum_{i\in N_{k}(x)}\left\|\begin{pmatrix}1\\ \frac{X_{i}-x}{\tau_{n,k}(x)}\end{pmatrix}\right\|_{2}^{3}.$
@@
Figure S.6: Real data analysis – ROC curve of the nearest neighbor classifier (top) and random forest classifier (bottom) with no dimension reduction (black curve), dimension reduction following the LLO$(\lambda=0)$ procedure (red curve) and dimension reduction following the LLO$(\lambda>0)$ procedure (blue curve). In each case, the prediction exercise is carried out on the selected testing set.